# Design and Characterization of a Cryogenic Vacuum Chamber for Ion Trapping Experiments

### v1

@@ Abstract
Trapped ions represent a widely used platform for quantum computing experiments. Fault-tolerant fidelities have been demonstrated for trapped-ion state preparation and readoutMyerson et al. (2008a); Sotirova et al. (2024); An et al. (2022), single-qubit gatesBrown et al. (2011); Smith et al. (2024), and two-qubit gatesBenhelm et al. (2008); Clark et al. (2021); Löschnauer et al. (2024); Moses et al. (2023). These operations have been combined into more complex algorithms within an architecture in which ions are shuttled among various locations between gate operations Pino et al. (2021); Moses et al. (2023); Srinivas et al. (2021). Preserving qubit coherence during long quantum algorithms requires isolation of the ion trap from external sources of noise such as mechanical cold-head vibrations and ambient magnetic field fluctuations. Achieving the highest possible detection fidelity requires high-efficiency light collection and detection, often constrained by the geometry of the ion trap and vacuum chamber. Here we present a cryogenic vacuum chamber designed with vibration isolation, in-vacuum magnetic shielding, a movable in-vacuum high numerical aperture (NA) objective, and an in-vacuum radio-frequency (rf) antenna for driving magnetic-dipole transitions in trapped ions confined to a surface-electrode trap. We describe vibrometer characterization of the trap vibrations, analysis of the ion-fluorescence collection efficiency and state-detection fidelity, and measurements of the coherence time of a magnetically sensitive qubit.
@@
The vacuum chamber is cooled using a two-stage Gifford-McMahon cryocooler (SHI Cryogenics RDK-415D2), powered by a helium compressor (Sumitomo F-70L), with 45 W of cooling power at the 50 K stage and 1.5 W at the 4.2 K stage. The expansion and compression cycle of the Gifford-McMahon cryocooler can create acoustic vibrations of greater than 20 mTomaru et al. (2004) that can couple to mechanical resonances in the chamber and displace the ion trap relative to external optics. The mechanical coupling between the cold head and chamber can be greatly suppressed by suspending the cold head on a separate mechanical support and employing helium to conduct heat between the chamber and cold head in a gas exchange space; flexible bellows seal gas inside the exchange space with reduced vibration propagation to the chamberDubielzig et al. (2021). Even with these isolation mechanisms in place, vibrations from the cold head can induce relative displacements up to 50-100 nm root-mean-square (RMS), resulting in fluctuations in optical phase and intensity at the ion trap that contribute to considerable errors in laser-based gatesPagano et al. (2018). These systems also experience slow drifts in position due to the support structure’s long lever arm in combination with slow variations in temperature within the chamber. Newly designed cryogenic systems further suppress mechanical oscillations by thermally anchoring the ion trap to the 4 K stage via flexible copper braidsMiossec et al. (2022a); Micke et al. (2019). Vibrations of less than 20 nm RMS were achieved by utilizing a 1.7 m long inertial pendulum to decouple the ion trap from the cryocooler in a separate mechanical roomMicke et al. (2019). We employ similar thermal anchoring to achieve low vibration levels but with a more compact design in which the ion trap is mechanically referenced to a floating optical table via short, thermally insulating support posts.
@@
A diagram of the cryogenic vacuum chamber is shown in Figure 1(a); the mechanical support structure is shown in further detail in Figure 2(a). The vacuum chamber contains two concentric copper radiation shields which are thermally anchored to the 4 K and 50 K stages with flexible copper braids. The ion trap is mounted inside the cold shield near 4 K; this shield is supported by the intermediate shield near 50 K via Vespel posts. The intermediate shield is then supported directly by the room-temperature, stainless steel chamber via Macor posts. The chamber is bolted to a floating optical table. In this configuration, the ion trap is mechanically referenced via short, stiff connections to the plane of the optical table and also therefore to all beam optics, including the lens stack (mounted below the table) used for ion imaging. The low thermal conductivity of both Vespel and Macor allow the cold and intermediate shielding assemblies to be sufficiently thermally isolated from each other and from the room-temperature vacuum chamberWoodcraft and Gray (2009). When the cold head is cooled to 5 K, the cold shield reaches 9.9 K and the intermediate shield reaches 85.9 K as measured by silicon diode sensors mounted inside the chamber.
@@
We characterize relative lateral displacements between the optical table and the cold shield with a commercial laser Doppler vibrometer (Polytec OFV-5000 with OFV-552 fiber-optic sensor heads). We secure the sensor head to the optical table and reflect the light off a 0.5-inch mirror mounted on the cold shield. We perform displacement measurements relative to the sensor head over intervals of 20 s and pass the signals through a 30 mHz high-pass filter to remove slow drifts [Fig. 2 (b)]. With the cold head powered off, we measure RMS vibrations of the cold shield to be 3.78(3) nm. We perform spectral analysis of the vibrations by performing a discrete Fourier analysis of the 20 s time series data with 0.05 Hz resolution bandwidth [Fig. 2 (c)]. With the cold head powered on, these vibrations increase to 7.61(4) nm. There is also a prominent peak at 1.2 Hz, the cryocooler pulsing frequency, and its harmonics. Vibrations at 30-300 Hz can be attributed to mechanical resonances of the floor and optical table which are excited when the cryocooler is on. At the achieved vibration amplitudes, we expect gate errors due to fluctuations in optical phase and intensity to be greatly reduced.
@@
State determination for trapped ions usually relies on collecting state-dependent fluorescence. Traditionally, the fluorescence is collected by an objective lens outside of the vacuum chamber and is directed to a camera or photomultiplier tube (PMT) for photon counting; with , detection error rates as low as in detection intervals as short as 145 sMyerson et al. (2008b); Burrell et al. (2010) have been achieved. The photon-detection efficiency of trapped-ion systems is often limited by the light-collection solid angle of the imaging objective. Recent efforts to improve readout fidelity and reduce detection durations focus on increasing the solid angle of light collection, either by moving an objective lens into the vacuum chamberCarter et al. (2024); Clark et al. (2014) or by integrating micro-fabricated optics and sensors directly into the trap structureTrue Merrill et al. (2011); Todaro et al. (2021); Reens et al. (2022). In our cryogenic system, we employ an in-vacuum objective with a working distance of 20.5 mm to achieve a numerical-aperture of 0.6, allowing us to collect 9.2% of the light scattered by the ion (10% from the solid angle with 91.7% transmission through the lens coating). The lens is mounted on a piezo-driven hexapod [see Fig. 1(a)], which allows us to remotely control the position and orientation of the optic with six degrees of freedom. Accounting for the losses from all optical interfaces in the imaging system and the quantum efficiency of the PMT, we predict an overall detection efficiency of 1.96% [Fig. 3(a)].
@@
To determine the true detection efficiency, we perform single-photon detection experiments with a trapped ion, similar to those described in Refs. Ghadimi et al., 2017; Clark et al., 2014. Relevant transition frequencies are shown in Figure 3(b). First, the ion is Doppler cooled for 5 s with 397 nm and 866 nm beams. Then, the 866 nm re-pump beam is turned off for 5 s, allowing the atom to decay to the level (lifetime 1 s). The 397 nm beam is turned off and a 5 s delay is inserted to ensure the 397 nm beam is fully extinguished. Finally, the 866 nm beam is turned on to pump population into the level (lifetime 7 ns). From here, the atom quickly decays to the level, generating a single 397 nm photon. This procedure is repeated for experimental trials, with the PMT detecting a photon on 1,770 occasions, yielding a detection efficiency of 1.77(4)% in good agreement with the predicted value.
@@
Higher collection efficiency allows us to perform faster detection, leading to reduced measurement errors from spontaneous decay. To determine the detection fidelity of our system, we perform experiments to compare fluorescence histograms measured from ions prepared in a "bright" state ( level) against those measured from ions prepared in a "dark" state ( level). We infer the ion state from photon-count thresholding. For a specified count threshold, above-threshold events in the dark histogram contribute to false-positive errors () and below-threshold events in the bright histogram contribute to false-negative errors (). The average error is calculated as . To determine the detection fidelity experimentally, we count the photons detected when an ion is prepared in the bright state or in the dark state during a 50s duration for trials each of bright and dark. To reduce the contributions of additional state-preparation errors to this detection fidelity estimate, we use a sequence of six laser pulses of 729 nm light to shelve the electronic population from the level to several states in the level for dark-state preparations, as shown in Fig. 3(c). The resulting detection histograms are plotted in Fig. 3(d). With a threshold of 5 or more photons used to indicate a bright ion, we calculate an average detection error of 3.7(4).
@@
Fluctuations of the magnetic field during quantum operations can lead to phase decoherence of magnetic-field sensitive qubit states. In Refs. Clark et al., 2021 and Hilder et al., 2022, magnetic field noise was the dominant noise source limiting operation fidelity. In compact room-temperature ion-trapping systems, external magnetic shielding has been employed to reduce magnetic field fluctuations near the ions and to improve the coherence of magnetic-field-sensitive qubits to greater than 1 sRuster et al. (2016). For a large cryogenic system, enclosing the entire vacuum chamber in multiple layers of magnetic shielding is far more difficult and costly. Instead, we designed and installed multi-layered magnetic shielding inside the vacuum chamber [Fig. 1(a) (5)]. Two layers of mu-metal are mounted to the inner surface of the vacuum chamber and thermalized to room temperature. An additional layer of Cryo-NeticMagnetic Shielding Corporation is installed inside the cold shield. Apertures in the shielding are required to allow optical access to the trap, but these openings can degrade the shielding performance. We mitigate this effect for the largest (imaging) aperture (under the trap) by including additional shielding layers around the in-vacuum objective and hexapod. Using COMSOL, we optimize the magnetic shielding geometry and predict a shielding factor in the central trapping region of greater than for fields applied vertically and greater than for fields applied laterally.
@@
To determine the efficacy of our magnetic shielding, we characterize the Ramsey coherence of the ground-state qubit with a magnetic sensitivity of 2.8 MHz/G. Fig. 4 shows the Ramsey contrast as the delay between /2-pulses is varied. The data exhibits a Gaussian decay with coherence times of 24(2) ms without a Hahn echo and 0.25(1) s with the echoHahn (1950). This is two orders of magnitude longer than the 3.8(4) ms Ramsey Hahn-echo coherence time measured in a similar cryogenic system in our lab which does not contain magnetic shielding.
@@
Dynamical decoupling is a powerful tool that can extend the coherence time of a qubit beyond the usual limitsMorong et al. (2023); Valahu et al. (2022). The ability to drive dynamical decoupling pulses simultaneously on many ions at disparate locations would be advantageous for many architectures. To this end, we design a critically coupled coil resonant near the ground state qubit transition frequency of 21.59 MHz with a quality factor of 20. Once installed within the magnetic shielding environment, however, the quality factor decreases to 5, as the nearby electrically lossy high-permittivity material couples to the circuit. Despite the ensuing reduction in rf magnetic field, we can perform qubit rotations with a -time of 40 s when driving the coil with 2 W. To test the performance of dynamical decoupling using the rf coil, we perform additional Ramsey coherence experiments with XY4 and XY32 (eight cycles of XY4) dynamical decoupling sequencesMaudsley (1986) (Fig. 4). We find the coherence times are extended to 0.72(2) s and 0.81(3) s with 4 and 32 re-focusing pulses, respectively.
@@
If we wish to perform dynamical decoupling pulses simultaneously at disparate trap locations, we require both a uniform biasing magnetic field and a uniform coil drive field. Initial COMSOL simulations indicate that the magnetic field produced by the coil should vary by less than 1% over the surface of the trap. Due to the presence of magnetic shielding, we require the magnets establishing our bias field to be installed within the inner shielding layer [Fig. 1(b) (9)]. Limited by available space within the cold shield, we do not achieve an optimally uniform magnet configuration. We measure a magnetic field gradient of approximately 350 G/m, such that ground state Zeeman splittings differ by 0.98 kHz/m. In our current configuration, the resulting pulse frequency and amplitude errors prevent us from performing global dynamical decoupling. In the future, improvements to the magnet configuration can reduce this non-uniformity, and some multi-pulse dynamical decoupling sequences can be implemented which are more robust to pulse amplitude errorsShappert et al. (2013).
@@
In this paper, we have presented the design and characterization of a cryogenic vacuum chamber designed to reduce error rates in trapped-ion experiments. We measure RMS vibrations of the cold shield to be 7.61(4) nm. Our experimental photon detection efficiency is 1.77%, allowing for fast readout of ion states. We achieve single-shot measurement fidelity of 99.9963(4)% with a detection interval of 50 s. We also evaluate our magnetic shielding performance by measuring Ramsey coherence times of a magnetic field sensitive qubit to be 24(2) ms, which extends to 0.25(1) s with a single Hahn-echo re-focusing pulse. Lastly, we utilize a radio frequency antenna to demonstrate XY4 and XY32 dynamical decoupling sequences which extend the coherence of our ion to 0.72(2) s and 0.81(3) s, respectively. We expect these improvements to our chamber design to greatly impact our achievable algorithmic fidelity.
@@
This work was done in collaboration with Los Alamos National Laboratory. The data that support the findings of this study are available from the corresponding author upon reasonable request. The authors have no conflicts to disclose.

### v2

@@ Abstract
Trapped ions represent a widely used platform for quantum computing experiments. Fault-tolerant fidelities have been demonstrated for trapped-ion state preparation and readoutMyerson et al. (2008); Sotirova et al. (2024); An et al. (2022), single-qubit gatesBrown et al. (2011); Smith et al. (2024), and two-qubit gatesBenhelm et al. (2008); Clark et al. (2021); Löschnauer et al. (2024); Moses et al. (2023). These operations have been combined into more complex algorithms within an architecture in which ions are shuttled among various locations between gate operations Pino et al. (2021); Moses et al. (2023); Srinivas et al. (2021). Preserving qubit coherence during long quantum algorithms requires isolation of the ion trap from external sources of noise. Mechanical cold-head vibrations can lead to gate amplitude errors when the ion vibrates with respect to the control optics. Ambient magnetic field fluctuations lead to dephasing of magnetic field sensitive states. Achieving the highest possible detection fidelity requires high-efficiency light collection and detection, often constrained by the geometry of the ion trap and vacuum chamber. Here we present a cryogenic vacuum chamber designed with vibration isolation, in-vacuum magnetic shielding, a movable in-vacuum high numerical aperture (NA) objective, and an in-vacuum radio-frequency (rf) antenna for driving magnetic-dipole transitions in trapped ions confined to a surface-electrode trap. We describe a vibrometer characterization of the trap vibrations, an analysis of the ion-fluorescence collection efficiency and state-detection fidelity, and measurements of the coherence time of a magnetically sensitive qubit.
@@
The vacuum chamber is pumped to Torr using a turbo pump, then an ion pump and the cryostat are both turned on. Once cold, the chamber achieves a base pressure of approximately Torr, as measured at the vacuum feedthroughs near the top of Figure 1(a). The chamber is cooled using a two-stage Gifford-McMahon cryocooler (SHI Cryogenics RDK-415D2), powered by a helium compressor (Sumitomo F-70L), with 45 W of cooling power at the 50 K stage and 1.5 W at the 4.2 K stage. The expansion and compression cycle of the Gifford-McMahon cryocooler can create acoustic vibrations of greater than 20 mTomaru et al. (2004) that can couple to mechanical resonances in the chamber and displace the ion trap relative to external optics. The mechanical coupling between the cold head and chamber can be greatly suppressed by suspending the cold head on a separate mechanical support and employing helium to conduct heat between the chamber and cold head in a gas exchange space; flexible bellows seal gas inside the exchange space with reduced vibration propagation to the chamberDubielzig et al. (2021). The Gifford-McMahon cryocooler and closed cycle exchange gas system are part of an ultra-low vibration assembly supplied by Lake Shore Cryotronics (SHI-4XG-UHV-15). Even with these isolation mechanisms in place, vibrations from the cold head can induce relative displacements up to 50-100 nm root-mean-square (RMS), resulting in fluctuations in optical phase and intensity at the ion trap that contribute to considerable errors in laser-based gatesPagano et al. (2018). These systems also experience slow drifts in position due to the support structure’s long lever arm in combination with slow variations in temperature within the chamber. Newly designed cryogenic systems further suppress mechanical oscillations by thermally anchoring the ion trap to the 4 K stage via flexible copper braidsMiossec et al. (2022); Micke et al. (2019). Vibrations of less than 20 nm RMS were achieved by utilizing a 1.7 m long inertial pendulum to decouple the ion trap from the cryocooler in a separate mechanical roomMicke et al. (2019). We employ similar thermal anchoring to achieve low vibration levels but with a more compact design in which the ion trap is mechanically referenced to a floating optical table via short, thermally insulating support posts.
@@
A diagram of the cryogenic vacuum chamber is shown in Figure 1(a); the mechanical support structure is shown in further detail in Figure 2(a). The vacuum chamber contains two concentric copper radiation shields which are thermally anchored to the 4 and 50 K stages with flexible copper braids. The ion trap is mounted inside the cold shield near 4 K; this shield is supported by the intermediate shield near 50 K via Vespel posts. The intermediate shield is then supported directly by the room-temperature, stainless steel chamber via Macor posts. The chamber is bolted to a floating optical table. In this configuration, the ion trap is mechanically referenced via short, stiff connections to the plane of the optical table and also therefore to all beam optics, including the lens stack (mounted below the table) used for ion imaging. The low thermal conductivity of both Vespel and Macor allow the cold and intermediate shielding assemblies to be sufficiently thermally isolated from each other and from the room-temperature vacuum chamberWoodcraft and Gray (2009). Thermal connections between the trap and cold shield assembly occur via two primary paths. First, the trap die is bonded to a ceramic pin grid array (CPGA), with 100 gold pins providing both electrical and thermal connections to a PCB filter board well thermalized to the surrounding cold shield. A second thermalization path is provided by a copper bridge mounted above the trap surface as shown in Figure 1(b). This component makes contact with the top surface of the alumina trap package and is also bolted to the cold shield via openings in the PCB and magnetic shielding. With the cold head at 5 K, a thermometer inside the cold shield registers a temperature of 9.9 K. This differential is due to 2.9 W of heat load flowing through the copper braids thermalizing the cold shield with conductance of 0.6 W/K. This load is dominated by 1.8 W conducted through 102 BeCu wires used for trap control, 0.4 W from the Vespel supports, and an estimated 0.6 W of rf dissipation.
@@
We characterize relative lateral displacements between the optical table and the cold shield with a commercial laser Doppler vibrometer (Polytec OFV-5000 with OFV-552 fiber-optic sensor heads). We secure the sensor head to the optical table and reflect the light off a 0.5 in mirror mounted on the cold shield. We perform displacement measurements relative to the sensor head over intervals of 20 s and pass the signals through a 30 mHz high-pass filter to remove slow drifts [Fig. 2 (b)]. With the cold head powered off, we measure RMS vibrations of the cold shield to be 3.78(3) nm. We perform spectral analysis of the vibrations by performing a discrete Fourier analysis of the 20 s time series data with 0.05 Hz resolution bandwidth [Fig. 2 (c)]. With the cold head powered on, these vibrations increase to 7.61(4) nm. There is also a prominent peak at 1.2 Hz, the cryocooler pulsing frequency, and its harmonics. Vibrations at 30-300 Hz can be attributed to mechanical resonances of the floor and optical table which are excited when the cryocooler is on. At the achieved vibration amplitudes, we expect gate errors due to fluctuations in optical phase and intensity to be greatly reduced.
@@
State determination for trapped ions usually relies on collecting state-dependent fluorescence. Traditionally, the fluorescence is collected by an objective lens outside of the vacuum chamber and is directed to a camera or photomultiplier tube (PMT) for photon counting; with , detection error rates as low as in detection intervals as short as 145 sMyerson et al. (2008); Burrell et al. (2010) have been achieved. The photon-detection efficiency of trapped-ion systems is often limited by the light-collection solid angle of the imaging objective. Recent efforts to improve readout fidelity and reduce detection durations focus on increasing the solid angle of light collection, either by moving an objective lens into the vacuum chamberCarter et al. (2024); Clark et al. (2014) or by integrating micro-fabricated optics and sensors directly into the trap structureTrue Merrill et al. (2011); Todaro et al. (2021); Reens et al. (2022); Knollmann et al. (2025). In our cryogenic system, we employ an in-vacuum objective with a working distance of 20.5 mm to achieve a numerical-aperture of 0.6, allowing us to collect 9.2% of the light scattered by the ion (10% from the solid angle with 91.7% transmission through the lens coating). The lens is mounted on a piezo-driven hexapod [see Fig. 1(a)], which allows us to remotely control the position and orientation of the optic with six degrees of freedom. Taking into account the losses from all optical interfaces in the imaging system and the quantum efficiency of the PMT, we predict an overall detection efficiency of 1.96% [Fig. 3(a)].
@@
To determine the true detection efficiency, we perform single-photon detection experiments with a trapped ion, similar to those described in Refs. Clark et al., 2014; Ghadimi et al., 2017. Relevant transition frequencies are shown in Figure 3(b). First, the ion is Doppler cooled for 5 s with 397 and 866 nm beams. Then, the 866 nm re-pump beam is turned off for 5 s, allowing the atom to decay to the level (lifetime 1 s). The 397 nm beam is turned off and a 5 s delay is inserted to ensure the 397 nm beam is fully extinguished. Finally, the 866 nm beam is turned on to pump population into the level (lifetime 7 ns). From here, the atom quickly decays to the level, generating a single 397 nm photon. This procedure is repeated for experimental trials, with the PMT detecting a photon on 1770 occasions, yielding a detection efficiency of 1.77(4)% in good agreement with the predicted value.
@@
Higher collection efficiency allows us to perform faster detection, leading to reduced measurement errors from spontaneous decay. To determine the detection fidelity of our system, we perform experiments to compare fluorescence histograms measured from ions prepared in a "bright" state ( level) against those measured from ions prepared in a "dark" state ( level). We infer the ion state from photon-count thresholding. For a specified count threshold, above-threshold events in the dark histogram contribute to false-positive errors () and below-threshold events in the bright histogram contribute to false-negative errors (). The average error is calculated as . To determine the detection fidelity experimentally, we count the photons detected when an ion is prepared in the bright state or in the dark state during a 50s duration for trials each of bright and dark. To reduce the contributions of additional state-preparation errors to this detection fidelity estimate, we use a sequence of six laser pulses of 729 nm light to shelve the electronic population from the level to several states in the level for dark-state preparations, as shown in Fig. 3(c). The resulting detection histograms are plotted in Fig. 3(d). With a threshold of 5 or more photons used to indicate a bright ion, we calculate an average detection error of 3.7(4). Note that the true detection fidelity of the ground-state qubit will be reduced from this value as an imperfect 729 nm transfer pulse to the optical qubit is first required. Techniques such as rapid adiabatic passage can be used to increase the fidelity of this transfer pulseWunderlich et al. (2007).
@@
Fluctuations of the magnetic field during quantum operations can lead to phase decoherence of magnetic-field sensitive qubit states. In Refs. Clark et al., 2021 and Hilder et al., 2022, magnetic field noise was the dominant noise source limiting operation fidelity. In compact room-temperature ion-trapping systems, external magnetic shielding has been employed to reduce magnetic field fluctuations near the ions and to improve the coherence of magnetic-field-sensitive qubits to greater than 1 sRuster et al. (2016). For a large cryogenic system, enclosing the entire vacuum chamber in multiple layers of magnetic shielding is far more difficult and costly. Instead, we designed and installed multi-layered magnetic shielding inside the vacuum chamber [Fig. 1(a) (5)]. Two layers of mu-metal are mounted to the inner surface of the vacuum chamber and thermalized to room temperature. Cryo-NeticMagnetic Shielding Corporation material lines the interior of the cold shield in both radial and axial directions with openings to accommodate cross-trap laser beams, fluorescence collection, and neutral calcium flux. Any openings can degrade the shielding performanceYashchuk et al. (2013), and we mitigate this effect for the largest (imaging) aperture (under the trap) by including additional shielding layers around the in-vacuum objective and hexapod. Using COMSOL, we optimize the magnetic shielding geometry and predict a shielding factor in the central trapping region of greater than for fields applied vertically and greater than for fields applied laterally.
@@
To determine the efficacy of our magnetic shielding, we characterize the Ramsey coherence of the ground-state qubit with a magnetic sensitivity of 2.8 MHz/G. Fig. 4 shows the Ramsey contrast as the delay between /2-pulses is varied. The data exhibit a Gaussian decay with coherence times of 24(2) ms without a Hahn echo and 0.25(1) s with the echoHahn (1950). This is two orders of magnitude longer than the 3.8(4) ms Ramsey Hahn-echo coherence time measured in a similar cryogenic system in our lab which does not contain magnetic shielding.
@@
Dynamical decoupling is a powerful tool that can extend the coherence time of a qubit beyond the usual limitsMorong et al. (2023); Valahu et al. (2022). The ability to drive dynamical decoupling pulses simultaneously on many ions at disparate locations would be advantageous for many architecturesYeh et al. (2023); Martínez-Lahuerta_2024. To this end, we design a critically coupled coil resonant near the ground state qubit transition frequency of 21.59 MHz with a quality factor of 20. Once installed within the magnetic shielding environment, the quality factor decreases to 5, as the nearby electrically lossy high-permittivity material couples to the circuit. Despite the ensuing reduction in the rf magnetic field, we can perform qubit rotations with a -time of 40 s when driving the coil with 2 W. To test the performance of dynamical decoupling using the rf coil, we perform additional Ramsey coherence experiments with XY4 and XY32 (eight cycles of XY4) dynamical decoupling sequencesMaudsley (1986) (Fig. 4). We find the coherence times are extended to 0.72(2) s and 0.81(3) s with 4 and 32 re-focusing pulses, respectively.
@@
If we wish to perform dynamical decoupling pulses simultaneously at disparate trap locations, we require both a uniform biasing magnetic field and a uniform coil drive field. Initial COMSOL simulations indicate that the magnetic field produced by the coil should vary by less than 1% over the surface of the trap. Due to the presence of magnetic shielding, we require the magnets establishing our bias field to be installed within the inner shielding layer [Fig. 1(b) (9)]. Limited by available space within the cold shield, we do not achieve an optimally uniform magnet configuration. We measure a magnetic field gradient of approximately 35 mT/m, such that ground state Zeeman splittings differ by 0.98 kHz/m. In our current configuration, the resulting pulse frequency and amplitude errors prevent us from performing global dynamical decoupling. In the future, improvements to the magnet configuration can reduce this non-uniformity, and some multi-pulse dynamical decoupling sequences can be implemented which are more robust to pulse amplitude errorsShappert et al. (2013).
@@
In this paper, we have presented the design and characterization of a cryogenic vacuum chamber designed to reduce error rates in trapped-ion experiments. We measure RMS vibrations of the cold shield to be 7.61(4) nm. Our experimental photon detection efficiency is 1.77%, allowing for fast readout of ion states. We achieve single-shot measurement fidelity of 99.9963(4)% with a detection interval of 50 s. We also evaluate our magnetic shielding performance by measuring Ramsey coherence times of a magnetic field sensitive qubit to be 24(2) ms, which extends to 0.25(1) s with a single Hahn-echo re-focusing pulse. Finally, we utilize a radio frequency antenna to demonstrate XY4 and XY32 dynamical decoupling sequences which extend the coherence of our ion to 0.72(2) s and 0.81(3) s, respectively. We expect these improvements to our chamber design to greatly impact our achievable algorithmic fidelity.
@@
This work was done in collaboration with Los Alamos National Laboratory. The authors have no conflicts to disclose. The data that support the findings of this study are available from the corresponding author upon reasonable request.